Mine is actually "MULTiPLyiNG POLyNOMiALs by MONOiALs".

I just don't get this at all.
the problem:
2(2m-3)-3(2m-1)=9
heres another one:
4(3a-1)-5(1-a)=8

can you please answer them=]

it will help me out alot..

Answer by checkley75(3666)

2(2M-3)-3(2M-1)=9
4M-6-6M+3=9
-2M=9+6-3
-2M=12
M=12/-2
M=-6 ANSWER.
4(3A-1)-5(1-A)=8
12A-4-5+A=8
13A=8+5+4
13A=17
A=17/13 ANSWER.